A labourer built 35m of stone wall. If she had built 2m less each day, it would have taken her 2 days longer. How many metres did she build each day, working at her usual rate?
Answer by Boreal(15235)

amount built per day is x m
time it took was 35/x days
if x-2 m, then it would have taken (35/x)+2 days longer
Therefore, (35/x-2)=(35/x)+2
Multiply everything by x (x-2)
35x=35(x-2)+2x(x-2)
35x=35x-70+2x^2-4x
2x^2-4x-70=0
x^2-2x-35=0, dividing by 2
(x-7)(x+5)=0
x=-5,7, use only 7 days
That is 5 meters per day
If she completed 5 meters per day, it would have taken 2 days longer, or 7 days.